a piece of wood from a recently cut tree shows 20 decays per minute.a wooden piece of same size placed in a museueum shows 2 decays per minute.if half life of c14 is 5730 years.then age of wooden piece placed in museum is approximately?

Grade:12

1 Answers

Abhishek Kumar
9 years ago
\lambdaN’ = 1/30 dps (dps =disintegration per second)
\lambdaN = 1/3 dps
So, N’/N = 1/10
e-\lambdat =1/10
Since, 1/\lambda = half life = 5730 yr
-\lambdat = ln(1/10)
So,
t = ln(10) X 5730 yr